find the equation in the form of y=mx+b through (2,2) perpendicular to y=3x+4
Answer by Maths68(1474)

Given
Point (x, y)=(2,2)
Line:
y=3x+4
Compare above equation with the equation of line slope-intercept form
y=mx+b
y=(3)x+4
m=3 and b=4
Slope of the given line m = 3 and y-intercept = b = 4
Since required line is perpendicular, the multiplication of the slopes of both lines result in (-1), therefore the slope of the required line will be (-1/3)
================================================================
Now we have a point (2,2) and slope (-1/3) of the required line we can easily find the required line put these values in the equation of slope-intercept form to find the y-intercept of the required line
y=mx+b
2=(-1/3)(2)+b
2=-2/3+b
2+2/3=b
6+2/3=b
8/3=b
b=8/3
y-intercept of the required line =b=8/3
================================================================
Put the values of ‘m’ and ‘b’ in equation of the line slope-intercept form
y=mx+b
y=(-1/3)x+8/3
Above equation is the required equation of the line in slope-intercept form.